# Sliding Puzzles Gym: A Scalable Benchmark for State Representation in Visual Reinforcement Learning

### v4

@@ Acknowledgements
The authors gratefully acknowledge the valuable insights and constructive discussions provided by Professors Marcos R. O. A. Maximo and Flávio H. T. Vieira.
@@ C.2.1 Analysis of Algorithmic Variants
Pretraining. As shown in Figure 12, pretraining provides clear benefits for PPO, especially with larger pools. While in-distribution pretraining provides strong gains, almost matching the performance of PPO with one-hot-based observations (see Section C.1.1), this represents an optimistic upper bound since real-world scenarios rarely permit task-specific pretraining. Interestingly, out-of-distribution pretraining also shows benefits compared to random initialization (90% vs 86% success at pool size 5), suggesting some transfer of useful visual features from the pretrained encoder. This indicates that even general-purpose pretrained encoders can provide a helpful initialization for RL tasks, though not matching the performance of task-specific pretraining.

### v5

@@ Acknowledgements
The authors gratefully acknowledge the valuable insights and constructive discussions provided by Professors Marcos R. O. A. Maximo and Flávio H. T. Vieira. We also thank the reviewers for their thoughtful feedback and constructive suggestions that helped improve this paper.
@@ C.2.1 Analysis of Algorithmic Variants
Pretraining. As shown in Figure 12, pretraining provides clear benefits for PPO, especially with larger pools. While in-distribution pretraining provides strong gains, almost matching the performance of PPO with one-hot-based observations (see Appendix C.1.1), this represents an optimistic upper bound since real-world scenarios rarely permit task-specific pretraining. Interestingly, out-of-distribution pretraining also shows benefits compared to random initialization (90% vs 86% success at pool size 5), suggesting some transfer of useful visual features from the pretrained encoder. This indicates that even general-purpose pretrained encoders can provide a helpful initialization for RL tasks, though not matching the performance of task-specific pretraining.